# Christian Ghiglino    Alastair Langtry

### v4

@@ 2.1 Machinery & Assumptions
The preferences in Eq (1) put our model into the class of linear-quadratic games on a network (see Jackson and Zenou (2015); Bramoullé and Kranton (2016) for reviews). As is standard in this class of games, some variant on Bonacich centrality will play an important role in equilibrium behaviour. And to ensure an equilibrium exists we will need the largest eigenvalue of the network to be strictly less than one (Ballester et al., 2006; Bramoullé et al., 2016).
@@
Definition 1 (Generalised Bonacich Centrality).
@@
$C^{b}_{j}=\sum_{k}\left[\sum_{t=0}^{\infty}H^{t}\ \right]_{jk}\cdot\frac{w_{k}}{\beta w_{k}+1}$ (2)
In the interest of brevity, we will just refer to this as ‘centrality’. It will also be convenient to define the average centrality of agents with a given identity. Formally, let $\bar{C}^{b}_{\theta}=\frac{1}{|I(\theta)|}\sum_{k\in I(\theta)}C^{b}_{k}$ be the average centrality of agents with identity $\theta$.
Note that our eigenvalue condition applies to the network $H$ (rather than $G$). It ensures that the infinite sum in Equation 2 converges, and hence that the centrality measure is well-defined.
@@
$|\lambda_{1}|<1$, where $\lambda_{1}$ is the largest eigenvector modulus of the matrix $H$.
Given the functional form we impose in Equation 1, in order to guarantee that status benefits are increasing in $Y(\theta_{j})$, we require that no agent’s centrality is ‘too far’ above the average. Formally, it amounts to the following (necessary and sufficient) restriction.
Assumption 2.
Assume that for all agents $j$:
$C^{b}_{j}<\frac{1}{\gamma}\cdot\frac{f(\bar{C}^{b}_{\theta_{j}})^{3}}{\gamma+\alpha f(\bar{C}^{b}_{\theta_{j}})},\ \text{ where }\ f(\bar{C}^{b}_{\theta_{j}})=\frac{\alpha}{4}\bar{C}^{b}_{\theta_{j}}+\frac{1}{4}\cdot(\alpha\bar{C}^{b}_{\theta_{j}})^{0.5}\cdot(\alpha\bar{C}^{b}_{\theta_{j}}+\frac{8\gamma}{\alpha})^{0.5}.$
Intuitively, each agent’s centrality must be less than some (increasing) function of the average centrality. It is clear that the higher the average centrality, the more dispersion is permitted.
We impose Assumptions 1 and 2 throughout the paper, and will not restate them for each result.
@@ 2.2 Discussion
Our utility formulation in Equation 1 closely follows Shayo (2009) – its two core components are the status benefits and dissonance costs. Shayo (2009), as in many other models, focuses on how people choose their identity. Therefore he treats the dissonance costs as depending only on the identity choice – not on actions. Given our different focus, we make the opposite choice: agents’ reference points – the ‘benchmark for normal’ that they want to conform to – is determined by people’s endogenous consumption choices. Note that in the social identity literature, the ‘benchmark for normal’ is typically called a prototype (including in Shayo (2009, 2020)), rather than a ‘reference point’. [^10] The ‘reference point’ terminology is more common in the literature on social comparisons, and stems from that literature’s more explicit link to behavioural economics.
@@
Further, we allow the reference point to depend on an agent’s position in a network. This follows the now-common notion that people compare themselves to those they actually interact with (e.g. Immorlica et al. (2017); Bramoullé and Ghiglino (2022)). The trade-off is that we treat identity as exogenously fixed in order to keep the model tractable and ensure clean results.
@@
Note that in line with typical social identity models, agents dislike being away from the reference point here – either above or below. This takes the view that people want to conform. It is motivated by significant evidence from social psychology, especially regarding descriptive norms (Cialdini et al., 1990; Cialdini and Goldstein, 2004).
This contrasts with the standard assumption in the social comparisons literature that people want to get ahead – and so dislike being below the reference point, but enjoy being above it. Nevertheless, it is possible to replace the social identity style desire to conform with a social comparisons style desire to get ahead. This change does not alter the key insights or the qualitative features of equilibrium behaviour. We show this in the Online Appendix. [^11] To remain highly tractable, it requires a linear, rather than quadratic, cost of consumption.
Further, empirical evidence in favour of each of these two views of how agents feel about being above the reference point is weak. In contrast, the idea that agents dislike being below the reference point—an idea common to both social identity and social comparisons models—finds strong empirical support.
@@
Before showing how our model rationalises the stylised facts set out in the introduction, we first analyse the model on its own terms. We characterise equilibrium behaviour, and show how it depends critically on an agent’s own centrality, and the average centrality of her identity group. Then we examine how the primitives of the model impact centrality.
@@ 3 Results
We start by solving the model: that is providing an exact characterisation of how agents behave in equilibrium.
@@
$x_{j}^{*}=\frac{1}{2}\left(\alpha+\frac{\gamma}{f(\bar{C}^{b}_{\theta_{j}})}\right)C_{j}^{b}.$ (3)
Note that $f(\bar{C}^{b}_{\theta_{j}})$ is defined as in 2. Centrality is the key determinant of equilibrium consumption. As in other linear-quadratic games, higher own centrality increases an agent’s own consumption. This is because agents with high centrality are more connected to others who themselves have high consumption – giving them a high reference point. In turn this pushes them to consume more in an effort to conform to the ‘benchmark for normal’ set by the reference point.
There is also high status from belonging to a group where the other members high consumption (which is driven by them having high centrality). But status from the group is imperfectly substitutable with the status frown own consumption. So when an agent belongs to a high status group, she has a lower marginal benefit of own consumption – so she does less of it. The substitutability between the two sources of status is key. When an agent belongs to a high status group, she feels less need to create status for herself through her own consumption. The following result formalises this discussion.
@@
$x_{j}^{*}$ is strictly increasing in $C_{j}^{b}$, and strictly decreasing in $\bar{C}^{b}_{\theta_{j}}$, for all $j$.
This links equilibrium behaviour to centrality. The next natural question is to ask what drives centrality. High centrality in our model comes from a combination of strong connections to neighbours (especially to those who are themselves well connected), connections to people with high income, and having high income oneself.
@@
$C_{j}^{b}$ is increasing in $w_{\ell}$, and in $G_{\ell m}$, for all $j,\ell,m$. These results are strict if and only if $j$ and $\ell$ share an identity and there is a walk from $\ell$ to $j$.
@@
More and stronger links raise an agent’s reference point – the benchmark she tries to conform to. This follows from the definition of the reference point. So she raises her consumption in response. In turn, this then pushes up her neighbours’ reference points, which raises their consumption, and so on. The effect ripples out through the network. Higher income makes consumption cheaper for an agent (importantly, at the margin) – so she chooses to consume more directly. This then raises her neighbours’ reference points, and so ripples out through the network in much the same way as with stronger links. In both cases, the ripple effect is contained within the affected identity group – other identity groups are unaffected.
@@
This is because we assumed that the reference point – the ‘benchmark for normal’ that agents feel pressure to conform with – depends only on those with the same identity. It captures that idea that I only feel pressure to conform to ‘people like me’. Those with a different identity are seen as different, and so not conforming with them does not present any issues.
A consequence of this feature is that homophily has a particularly clean impact on centrality. Homophily is the tendency for people to be disproportionately connected to those similar to themselves, and is one of the most ubiquitous features of social networks (Jackson, 2008; Pin and Rogers, 2016). Higher homophily means more links between people with the same identity (at the expense of fewer cross-identity links). And because it is only the within-identity links that matter for our centrality measure, higher homophily has much the same impact as higher network density.
To show this properly, we now need a formal definition of homophily. There are many in the literature, but we use the homophily index as one of the simplest available. The homophily index is specific to an agent and simply measures the proportion of total links that are with agents in the same group. And we naturally take the identity groups as the groups of interest.
Definition 2 (Homophily index: from Currarini et al. (2009)).
For an agent $j$ in group $\theta$:
$\phi_{j}=\frac{\sum_{k\in I(\theta)}G_{jk}}{\sum_{k}G_{jk}}$ (4)
Additionally, we need to look at some particular type of changes to the network. We go for the simplest possible: swapping a link $G_{jk}$ for a link $G_{j\ell}$. [^12] It is clear that any change to the network structure that holds network density (i.e. the weighted total number of links) constant can be created from a series of link swaps. It is clear from 2 that changing a link $G_{jk}$ to a link $G_{j\ell}$ only affects the homophily index of agent $j$. So, when examining the effect of changing a link, we can conflate changes in the homophily index for agent $j$ and homophily more broadly without problems. It also follows easily from the definition that homophily falls when $j,k$ share an identity and $\ell$ has a different one, that homophily rises when $j,\ell$ share an identity and $k$ has a different one, and is left unchanged otherwise. [^13] To see this, first notice that the denominator on the right-hand side of eq. 4 never changes – this is because we are only changing links. So homophily falls if changing the link reduces the number of within-identity links. Similarly, homophily rises if changing the link increases the number of within-identity links. These other cases are that (a) $j,k$ and $\ell$ all share an identity, and (b) $j$ has a different identity to both $k$ and $\ell$. In case (b) it does not matter whether or not $k$ and $\ell$ have the same identity because the link between $k$ and $\ell$ (if it exists) is not affected.
In this setting, we can then say that changing links in a way that increases homophily will increase centralities. And similarly, changing links in a way that decreases homophily will decrease centralities.
Remark 3.
Swap a link $G_{jk}$ for a link $G_{j\ell}$.
(i) if homophily rises, then $C_{m}^{b}$ rises for all $m\in I(\theta_{j})$,
(ii) if homophily falls, then $C_{m}^{b}$ falls for all $m\in I(\theta_{j})$.
This result is driven by the fact that only within-identity links that matter for equilibrium behaviour (recall that it is the matrix $\widehat{G}$ rather than the matrix $G$ that appears in the definition of centrality). This is because it is only these links that drive the reference point. In our model, agents do not care per se about the decisions of those from different identity groups. So changing a link in the network $G$ in a way that increases homophily adds a links in the network $\widehat{G}$ – which is the one that enters into the definition of centrality. [^14] Note that this observation also means that swapping a between-identity link for a different between-identity link has no effect. Swapping two different within-identity links is more complicated. There, both adding the link $G_{j\ell}$ and removing $G_{jk}$ have an impact on the network $\widehat{G}$. However, high centrality is driven by being linked to people who themselves have high centrality. So $j$’s centrality rises if swapping the link connects her to someone who has higher centrality. Conversely, if the swap connects her to someone with lower centrality, her own centrality falls. This follows straightforwardly from Langtry (2022, Proposition 3).
The effect of discrimination.
In reality, the status of different identity groups will depend on a number of other (exogenous) factors in addition to the visible good consumption of the group’s members that this paper focuses on. [^15] It would be straightforward to add it to the model. Simply define $Y(\theta_{j})=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}x_{k}+E(\theta_{j})$, where $E(\theta_{j})>0$ is some exogenous parameter capturing status from other (exogenous factors). This exogenous status component may be linked to discrimination or marginalisation in a society. Groups that are discriminated against or marginalised would have a low exogenous status component, while favoured groups would have high exogenous status.
Importantly, this will affect their consumption choices. All else equal, agents belonging to the marginalised groups will be pushed towards greater visible good consumption—and away from non-visible good consumption. [^16] We do not model other goods explicitly, but the convex cost function captures the opportunity cost of foregoing consumption of non-visible goods in stylised manner. This is because they use visible good consumption to compensate for their otherwise low status. This follows from one of the key features of our model – that agents get status both from their identity group and from their own consumption, and, crucially, that these two sources are imperfect substitutes.
The broader impacts of this will depend on the properties of visible goods compared to non-visible ones. While what constitutes a visible (status generating) good will depend on context, the standard view of such goods is as luxury consumption goods (e.g. Charles et al. (2009)). Non-visible goods cover everything else, but importantly include durable goods and investments in capital. So discrimination and marginalisation pushed agents away from spending on non-visible goods, and by extension may push them away from investments in capital.
This suggests a new potential harm of discrimination: it pushes agents towards greater visible good consumption, at the loss of investment in non-visible capital goods. To our knowledge, this is a novel prediction about how discrimination and marginalisation affect behaviour. Formal analysis of welfare effects is beyond the scope of this paper – we would need to model non-visible goods explicitly and examine how agents benefit from them.
@@ 4 The effects of changes in income
In the introduction, we set out a series of stylised facts about how changes in incomes feed through into conspicuous consumption decisions. And we noted that existing models struggle to rationalise them all as a coherent whole. The main aim of this section is therefore to show how our model fits all of these facts simultaneously.
In the previous section we found that a change in incomes affects both agents’ centralities and the average centralities of their identity groups – and that these have opposing effects on equilibrium behaviour. While well-known intuitions regarding Bonacich centrality style measures suggest that the impact of a change to incomes (or other primitives) have a larger impact on agents closer in the network to the change. So the key idea is that for agents located ‘close’ to the change, the effect on own centrality will dominate the effect on average centrality. And that the reverse applies for agents located ‘far from’ the change. But it is difficult to state precise results for general network structures.
To show this clearly we will therefore focus on a stark benchmark setting where agents are split up into disconnected ‘communities’. This will allow us to separate each channel cleanly. And it gives us a stylised way of thinking about ‘near’ and ‘far’ in the network. A pair of agents who share a community are ‘near’ to one another in the network, and a pair in different communities are ‘far apart’.
@@
The network $\widehat{G}$ consists of $N$ strongly connected components, [^17] A walk in $\widehat{G}$ from $j$ to $k$ is a sequence of agents $j,j^{\prime},...,k^{\prime},k$ such that $\widehat{G}_{j^{\prime}k^{\prime}}>0$ for all $j^{\prime},k^{\prime}$ in the sequence. A strongly connected component is a set of agents such that for all agents $j,k$ in the set, there exists a walk from $j$ to $k$. with typical component $n$, and where $N$ is large. Call each component a community.
Changes in one community will only affect centralities within that one community, and will not spill over into other communities. This means a change in one community affects other communities only through the impact on average centrality. Further, assuming that there are many communities guarantees that a change in community $n$ has a larger impact on the centrality of the agents in that community than it does on the average centrality of all agents. In summary, following a change in community $n$, the ‘own centrality’ channel dominates within community $n$, and the ‘average centrality’ channel dominates in all other communities.
@@
    In the US [resp. India], people consume less visible goods when the average income of others of the same race in their state [others of the same caste/religious group in their district] is higher (Charles et al., 2009; Khamis et al., 2012).
@@
    In South Africa, people’s consumption of visible goods is more sensitive to changes in income of their reference group when they are richer (Kaus, 2013).
Note that we defer the fourth stylised fact to later in the section. This is because it relates to changes in inequality, and is less clear-cut – and so needs dedicated analysis. With these facts in mind, we can now state our result.
@@
(iii) For all agents $j$ in communities other than $n$, the marginal impact on $j$’s consumption is increasing in $j$’s own income.
@@
The mapping from 2 into the stylised facts is fairly direct. An increase in incomes in other communities corresponds to higher average income of others of the same race in the same state, as in Charles et al. (2009). [^18] We use the term ‘income’ in our model for convenience. Where appropriate, this should be taken to mean the ‘permanent income’ examined by Charles et al. (2009) in the empirical analysis. An increase in incomes within an agent’s own community corresponds to raising her neighbours’ incomes, as in Kuhn et al. (2011). Note that 2 applies only to agents who share an identity, but Kuhn et al. (2011) does not consider identity in their analysis. Nevertheless, strong homophily and sorting along geographic lines are well-documented phenomena within economics, sociology McPherson et al. (2001); Schelling (2006) and other disciplines. So it is very likely that at least some of an agent’s geographic neighbours do in fact share her identity. The third stylised fact is even more direct: an ‘increasing marginal impact’ is simply a formalisation of the ‘greater sensitivity’ found by Kaus (2013).
From a more technical standpoint, 2 follows from 1 when focusing attention on the communities network. In that case, an increase in the incomes of some agents with identity $\theta$ in community $n$ does two things. First, it increases the centralities of agents with identity $\theta$ in community $n$. Second, this then increases the average centrality of identity $\theta$. Importantly, it does not affect the centrality of identity $\theta$ agents in other communities. So consumption falls in the communities other than $n$ – because average centrality rose, with no impact on own centrality. And consumption rises in community $n$ – because the effect on own centrality dominates the effect on average centrality. The sensitivity result is also immediate: higher own income increases own centrality, and this acts multiplicatively, making consumption more sensitive to other changes.
@@
More importantly, 2 illuminates the mechanism driving this behaviour. An increase in incomes for some agents with identity $\theta$ in community $n$ raises their consumption directly. This then raises the reference points of their neighbours, raising the benchmark these neighbours are trying to conform to – and hence raising their consumption. This effect ripples out through the whole community (see discussion following 2). This then raises the average consumption of agents with identity $\theta$. So agents with that identity gain more status from belong to that identity group. So those in other communities then feel less need to seek status through their own consumption of visible goods.
But for agents in community $n$, the changes are concentrated ‘close’ to them in the network. So the pressure to conform to a higher benchmark is strong relative to the gains from their whole identity group having higher status. It is this pressure from a higher benchmark that pushes up consumption – and this is the force that dominates inside community $n$.
@@
While Propositions 1 and 2 are clear on how the third stylised fact – the greater sensitivity of agents with higher incomes – is captured mechanically, the headline intuition is relates to the way income affects the costs of consumption. In particular, agents with higher income have a lower marginal cost of consumption, and this marginal cost rises more slowly. When the marginal status benefits of consumption rise, as happens when $Y(\theta_{j})$ decreases, it should not come as a surprise that agents increase their consumption in response. But by how much they adjust their consumption depends on how quickly the marginal cost rises as consumption rises. It is precisely higher income agents whose marginal costs rise more slowly, and so who make greater adjustments.

### v7

@@ 2.1 Machinery & Assumptions
The preferences in Eq (1) put our model into the class of linear-quadratic games on a network (see Jackson and Zenou (2015); Bramoullé and Kranton (2016) for reviews). In particular, the quadratic dissonance term is similar in spirit to Ushchev and Zenou (2020). As is standard in this class of games, some variant on Bonacich centrality will play an important role in equilibrium behaviour.
@@
Definition 1 (Modified Bonacich Centrality).
@@
$C_{j}=\sum_{k}\left[(I-H)^{-1}\right]_{jk}\cdot\frac{w_{k}}{\beta w_{k}+1}$ (2)
In the interest of brevity, we will just refer to this as ‘centrality’. It will also be convenient to define the average centrality of agents with a given identity. Formally, let $\mathcal{C}_{\theta}=\frac{1}{|I(\theta)|}\sum_{k\in I(\theta)}C_{k}$ be the average centrality of agents with identity $\theta$. And to ensure centrality is well-defined (and that an equilibrium exists) impose the eigenvalue standard to linear-quadratic games (Ballester et al., 2006; Bramoullé et al., 2016). We also need the marginal benefit of consumption $S^{\prime}_{x}(\cdot)$ to be bounded.
@@
(i) $|\lambda_{1}|<1$, where $\lambda_{1}$ is the largest eigenvalue modulus of the matrix $G$,
(ii) $\rho<\infty$, where $\psi\equiv\max_{j}\{C_{j}\}<\infty$ and $\rho\equiv S^{\prime}_{x}(\frac{1}{\psi(1+\beta)})$.
Part (ii) ensures that consumption is finite for all agents. It also ensures that the ratio of average consumptions, $Y_{\theta}$ is also finite. Together, this allow us to place our assumptions on the function $S(x,Y)$ only for a finite range. Specifically, for $x\in[0,X]$, where $X\geq\psi\rho$ and $Y\in[\frac{1}{Z},Z]$ where $Z\geq\psi(1+\beta)$. [^11] This is important, because the assumptions we impose on $S(x,Y)$ cannot hold for an infinite range. Note that our characterisation of the range in which our assumptions $S^{\prime}_{x}>0,S^{\prime}_{Y}>0,S^{\prime\prime}_{xY}<0$ must hold could be tightened if we imposed some additional, but still mild, assumptions. We discuss this in Appendix B We impose Assumption 1 throughout the paper and will not restate it for each result.
@@ 2.2 Discussion
Before going further, it is helpful to briefly discuss three key features of the modelling. First, the theory of where status comes from. Second, that we have only one good. And third, the dissonance term: its motivation and relationship to existing work on social identity.
Status substitution.
Our model is one where consuming the status good brings utility directly, rather than being a signal for unobservable wealth. We discussed this in the introduction, and will not repeat ourselves here. Nevertheless, it is important to note that we need status good consumption to be visible. This is the same as in signalling models, but the reason is different. Here, agents must see their neighbours’ consumption in order to form their reference point.
Importantly, we assume (imperfect) substitutability between the direct benefits from own consumption and the status benefits derived from social identification. Such a trade-off between the self and the group identity is at the heart of social identity theory (Tajfel and Turner, 1979). There is extensive empirical evidence for substitution, in particular that individuals with low personal self-esteem or competence increase reliance on group identification. Seminal papers include (Abrams and Hogg, 1988; Hogg, 2000; Crocker and Luhtanen, 1990)) while recent empirical studies confirming this include (Cambon and Yzerbyt, 2018; Gonzales-Backen et al., 2015).
Single good.
To help present clean results, our model has only one good, and uses a quadratic cost term to capture the opportunity cost of foregoing other (unmodelled) goods. This cost term depends explicitly on income, as the level of income affects the marginal opportunity cost of other goods. This income term implicitly accounts for (exogenous) prices – as these also affect the marginal opportunity cost.
We can also view this setup as a second order approximation of a two-good model with exogenous prices and budget. Formally, there is a status good $x$, a composite (non-status) good $z$, and additively separable utility $u(x,z)=v(x)+s(z)$, where $v(\cdot)$ and $s(\cdot)$ are increasing, strictly increasing and concave. Let $p$ denote the relative price of the status good, $W$ denote income. Substituting in the (binding) budget constraint $px+z=W$ and then taking a second order Taylor approximation of $u(x,z)$ delivers utility as a quadratic function of $x$. Finally, notice that Equation 1 – while written in a particular way to capture notions of status and dissonance – is ultimately a quadratic function of $x$.
Note that our model – even viewed as an approximation to a two-good case – treats prices as fixed. It takes a partial equilibrium view that ignores the supply side of the economy. We do this because general equilibrium effects in social comparisons model are known to be extremely complex (Ghiglino and Goyal, 2010), and more importantly because they would not bring additional insights regarding the type of questions we address in this paper.
Dissonance.
@@
Our utility formulation in Equation 1 closely follows Shayo (2009) – in that its two core components are the status benefits and dissonance costs. Agents bear dissonance costs from departing from the ‘benchmark’ for their identity group. Here, this ‘benchmark’ is called a reference point. [^12] Note that in the social identity literature, the ‘benchmark’ is typically called a prototype (including in Shayo (2009, 2020)), rather than a ‘reference point’. The ‘reference point’ terminology is more common in the literature on social comparisons, and stems from that literature’s more explicit link to behavioural economics. Shayo (2009) focuses on how people choose their identity. So for tractability he treats the dissonance costs as depending only on the identity choice – not on consumption.
@@
In contrast, we assume dissonance costs depend on consumption, and that the reference point (the ‘benchmark’) is also determined by endogenous consumption choices. Further, we allow the reference point to depend on an agent’s exact position in a network. This follows the now-common notion that people compare themselves to those they actually interact with (e.g. Ghiglino and Goyal (2010); Immorlica et al. (2017); Bramoullé and Ghiglino (2022)). The trade-off is that we treat identity as exogenously fixed in order to keep the model tractable and ensure clean results.
A limitation of our reference point is that it only depends on neighbours with the same identity. Intuitively, this is a highly stylised way of capturing the view that an agent’s benchmark for normal – that they want to conform to – is set predominantly by those in their own identity group. More prosaically, it is important for tractability.
@@
Nevertheless, we follow the standard view in social identity models that people want to conform: they dislike being either above or below the reference point. This is motivated by significant evidence from social psychology, especially regarding descriptive norms (Cialdini et al., 1990; Cialdini and Goldstein, 2004). But it stands in contrast to the standard assumption in the social comparisons literature that people want to get ahead: they dislike being below the reference point. Ultimately, changing a social comparisons style desire to get ahead would not alter the key insights of our model. We show this in Appendix C.
@@
Before showing how our model rationalises the stylised facts set out in the introduction, we first analyse the model on its own terms. We characterise equilibrium behaviour, and show how it depends critically on an agent’s own centrality, and the average centralities of the identity groups. Then we examine how the primitives of the model impact centralities.
@@ 3 Results
We start by solving the model: that is, providing an exact characterisation of how agents behave in equilibrium.
@@
$x_{j}^{*}=S^{\prime}_{x}(Y_{\theta_{j}}^{*})\cdot C_{j},\text{ where }Y_{\theta_{j}}^{*}\text{ solves }Y_{\theta_{j}}=\frac{S^{\prime}_{x}(Y_{\theta_{j}})\mathcal{C}_{\theta_{j}}}{S^{\prime}_{x}(\frac{1}{Y_{\theta_{j}}})\mathcal{C}_{-\theta_{j}}}$ (3)
Centrality is the key determinant of equilibrium consumption. As in other linear-quadratic games, higher own centrality increases an agent’s consumption. Here, this is because agents with high centrality are more connected to others who themselves have high consumption – giving them a high reference point. In turn this pushes them to consume more in an effort to conform to the ‘benchmark for normal’ set by the reference point.
There is also high status from belonging to a group where the other members have high consumption (which is driven by them having high centrality). But status from the group is imperfectly substitutable with the status from own consumption. So when an agent belongs to a high status group, she has a lower marginal benefit of own consumption – and so she consumes less. The substitutability between the two sources of status is key. When an agent belongs to a high status group, she feels less need to create status for herself through her own consumption.
But group status is relative. High consumption in the other identity group – driven by high centrality in that group – reduces the group status. By identical logic, this raises consumption. The relative notion of status drives cross-group effects: what happens in one identity group affects behaviour in the other – even though we assumed away any direct benchmarking against one another through the reference point. The following result formalises this discussion.
@@
$x_{j}^{*}$ is strictly increasing in $C_{j}$ and in $\mathcal{C}_{-\theta_{j}}$, and strictly decreasing in $\mathcal{C}_{\theta_{j}}$, for all $j$.
This links equilibrium behaviour to centrality. However, note that the comparative static on $C_{j}$ is just a thought experiment. Changes to primitives of the model will change the centrality of multiple agents, so changing $C_{j}$ while holding all else constant is not possible in practice. The next natural question is to ask what drives centrality. Unsurprisingly, high centrality in our model comes from a combination of strong connections to neighbours (especially to those who are themselves well connected), connections to people with high income, and having high income oneself.
@@
$C_{j}$ is increasing in $w_{\ell}$, and in $G_{\ell m}$, for all $j,\ell,m$. These results are strict if and only if $j$ and $\ell$ share an identity and there is a walk from $\ell$ to $j$.
@@
More and stronger links raise an agent’s reference point – the benchmark she tries to conform to. So she raises her consumption in response. In turn, this then pushes up her neighbours’ reference points, which raises their consumption, and so on. The effect ripples out through the network. Similarly, higher income makes consumption cheaper for an agent (importantly, at the margin) – so she chooses to consume more directly. This then raises her neighbours’ reference points, and so ripples out through the network in much the same way as with stronger links. In both cases, the ripple effect is contained within the affected identity group – other identity groups are unaffected.
@@
The ’containment’ of the ripple effect is driven by the fact that agents only consider those who share their identity when setting their reference point. It captures the idea that I only feel pressure to conform to ‘people like me’. Those with a different identity are seen as different, and so not conforming with them does not present any issues.
So far, 1 links centralities to consumption decisions, and 2 links incomes and network structure to centralities. But the overall link between incomes and consumption can be ambiguous. When incomes within an agent’s own identity group change, there are two effects. First, an increase in someone else’s income increases her own centrality. And second, it increases the average centrality of her group. These push in different directions (1). In contrast, an increase in income among agents in the other identity group has a clear effect. It must increase $j$’s consumption. This is because it reduces her group status without any impact on her own centrality. The following result states this formally.
Proposition 2.
(i) if $\theta_{j}\neq\theta_{k}$, then $x_{j}^{*}$ is strictly increasing in $w_{k}$. (ii) if $\theta_{j}=\theta_{k}$, then
[…]
This shows that spillovers across identity groups are clear – higher income in one identity group always drives higher consumption in the other. And for changes within an identity group, it decomposes the two contrasting forces. This helps us to see when the own centrality effect will dominate – pushing consumption up – and when the average centrality effect will dominate – pushing consumption down. But it is difficult to say in general exactly what network structure will lead one effect versus the other to dominate.
Well-known intuitions regarding Bonacich centrality style measures suggest that a change to incomes (or other primitives) has a larger impact on agents closer in the network to the change. So the key idea is that for agents located ‘close’ to the change, the effect on own centrality will dominate the effect on average centrality. And that the reverse applies for agents located ‘far from’ the change.
When we turn to explaining the stylised facts in the next sections, we will focus on a stark benchmark setting where agents are split up into disconnected ‘communities’. This will allow us to separate each channel cleanly. And it gives us a precise, if stylised, way of thinking about ‘near’ and ‘far’ in the network. A pair of agents who share a community are ‘near’ to one another in the network, and a pair in different communities are ‘far apart’.
@@ 4 The effects of changes in income
In the introduction, we set out a series of stylised facts about how changes in incomes feed through into how people choose to consume status goods. And we noted that existing models struggle to rationalise them all as a coherent whole. The aim of this section is to show how our model fits the first four of these six facts simultaneously. It will use the results from Section 3. The next two sections will show how our model fits the fifth and sixth stylised facts respectively, as well as building some additional results related to them.
As noted at the end of Section 3, we consider a set of disconnected ‘communities’. This allows us to isolate the channels identified in 2.
@@
The network $\widehat{G}$ consists of $N$ strongly connected components, [^13] A walk in $\widehat{G}$ from $j$ to $k$ is a sequence of agents $j,j^{\prime},...,k^{\prime},k$ such that $\widehat{G}_{j^{\prime}k^{\prime}}>0$ for all $j^{\prime},k^{\prime}$ in the sequence. A strongly connected component is a set of agents such that for all agents $j,k$ in the set, there exists a walk from $j$ to $k$. with typical component $n$, and where $N$ is sufficiently large. All components contain the same number of agents, and there are an equal number of components of each identity. Call each component a community.
Changes in community $n$ will only affect centralities within that community, and will not spill over into other communities. This means a change in one community affects other communities only through the impact on average centrality. Further, assuming that there are many communities guarantees that a change in community $n$ has a larger impact on the centrality of the agents in that community than it does on the average centrality of all agents. In summary, following a change in community $n$, the ‘own centrality’ channel dominates within community $n$, and the ‘average centrality’ channel dominates in all other communities.
@@
    In the US [resp. India], people consume less status goods when the average income of others of the same race in their state [others of the same caste/religious group in their district] is higher (Charles et al., 2009; Khamis et al., 2012).
@@
    In India, low caste households’ consumption of status goods is higher when Upper Caste households are richer (Bellet and Colson-Sihra, 2018).
  4. 4.
    In South Africa, people’s consumption of status goods is more sensitive to changes in income of their reference group when they are richer (Kaus, 2013).
With these facts in mind, we can now state our result.
@@
(iii) Consumption increases for all agents with identity $\theta^{\prime}\neq\theta$.
@@
(iv) For all agents $j$ in communities other than $n$, the marginal impact on $j$’s consumption is increasing in $j$’s own income.
@@
The mapping from 3 into the stylised facts is fairly direct. An increase in incomes in other communities corresponds to higher average income of others of the same race in the same state, as in Charles et al. (2009). [^14] We use the term ‘income’ in our model for convenience. Where appropriate, this should be taken to mean the ‘permanent income’ examined by Charles et al. (2009) in their empirical analysis. An increase in incomes within an agent’s own community corresponds to raising her neighbours’ incomes, as in Kuhn et al. (2011). Note that 3(ii) applies only to agents who share an identity, but Kuhn et al. (2011) does not consider identity in their analysis. Nevertheless, strong homophily and sorting along geographic lines are well-documented phenomena within economics, sociology (McPherson et al., 2001; Schelling, 2006) and other disciplines. So it is very likely that at least some of an agent’s geographic neighbours do in fact share her identity.
The third stylised fact is even more direct: an ‘increasing marginal impact’ is simply a formalisation of the ‘greater sensitivity’ found by Kaus (2013). Finally, an increase in incomes of one identity group (Upper Caste households) leads to an increase in visible/status goods consumption of other identity groups (low caste households). This is the fourth stylised fact.
@@
From a more technical standpoint, 3 follows directly from 1 when focusing attention on the communities network. Importantly, 1 illuminates the mechanism that drives this behaviour. An increase in incomes for some agents with identity $\theta$ in community $n$ raises their consumption directly. This then raises the reference points of their neighbours, raising the benchmark these neighbours are trying to conform to – and hence raising their consumption (Fact 2). This effect ripples out through the whole community (see discussion following 2). This then raises the average consumption of agents with identity $\theta$. So agents with identity $\theta$ gain more status from belonging to that identity group. So those with identity $\theta$ in other communities feel less need to seek status through their own consumption of status goods (Fact 1). Conversely, agents in other identity groups now have lower status – because status is a relative concept – and so feel a greater need to consume status goods (Fact 4).
@@
But for agents in community $n$, the changes are concentrated ‘close’ to them in the network. So the pressure to conform to a higher benchmark is strong relative to the gains from their whole identity group having higher status. It is this pressure from a higher benchmark that pushes up consumption – and this is the force that dominates inside community $n$ (Fact 1). Finally, while Propositions 1 and 3 are clear on how the third stylised fact is captured mechanically, the headline intuition is related to the way income affects the costs of consumption. In particular, agents with higher income have a lower marginal cost of consumption, and this marginal cost rises more slowly. When the marginal status benefits of consumption rise, as happens when $Y_{\theta_{j}}$ decreases, it should not come as a surprise that agents increase their consumption in response. But by how much they adjust their consumption depends on how quickly the marginal cost rises as consumption rises. It is precisely higher income agents whose marginal costs rise more slowly, and so who make greater adjustments.